OpenStudy (anonymous):

HELP AM I RIGHT? Find the slope of a line passing through (6, 7) and (2, 9)(Points : 2) -2 -4 I PUT -1/2

OpenStudy (anonymous):

???

hartnn (hartnn):

yes, you are correct, -2/4 = -1/2 is the slope.

OpenStudy (anonymous):

slope=9-7/2-6 slope=2/-4=-1/2